A car manufacturer is interested in estimating the true mean fuel consumption in real-world city driving for one of its new m7
Answer #1

Solution :

Given that,

Point estimate = sample mean = \bar x = 9.4

sample standard deviation = s = 0.343

sample size = n = 10

Degrees of freedom = df = n - 1 = 10 - 1 = 9

a) At 95% confidence level

\alpha = 1 - 95%

\alpha =1 - 0.95 =0.05

\alpha/2 = 0.025

t\alpha/2,df = t0.025,9 = 2.262

Margin of error = E = t\alpha/2,df * (s /\sqrtn)

= 2.262 * ( 0.343 / \sqrt 10)

Margin of error = E = 0.245

b) At 90% confidence level

\alpha = 1 - 90%

\alpha =1 - 0.90 =0.10

\alpha/2 = 0.05

t\alpha/2,df = t0.05,9 = 1.833

Margin of error = E = t\alpha/2,df * (s /\sqrtn)

= 1.833 * ( 0.343 / \sqrt 10)

Margin of error = E = 0.199